# Emulation Capacity between Idempotent Channels

### v1

@@ 3 Decomposition properties of idempotent channels
Let $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$ and $\widetilde{\mathcal{H}}=\bigoplus_{k=1}^{K}\widetilde{\mathcal{H}}_{k,1}\otimes\widetilde{\mathcal{H}}_{k,2}$ be such that
[…]
Then, let
$\mathcal{E}:x\in\mathcal{L}(\mathcal{H})\mapsto\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k}\text{ and }\mathcal{D}:x\in\mathcal{L}(\widetilde{\mathcal{H}})\mapsto\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(\widetilde{P}_{k}x\widetilde{P}_{k})U_{k}\otimes\rho_{k}$
with a set of states $\{\widetilde{\omega}_{k}\in\mathcal{D}(\widetilde{\mathcal{H}}_{k,2})~|~k\in[K]\}$ and, for all $k\in[K]$, a unitary $U_{k}:\mathcal{H}_{k,1}\mapsto\widetilde{\mathcal{H}}_{k,1}$. These unitaries allow to identify $\mathcal{L}(\mathcal{H}_{k,1})$ with $\mathcal{L}(\widetilde{\mathcal{H}}_{k,1})$ for each $k\in[K]$. Then, for $x\in\mathcal{L}(\mathcal{H})$, we have for any $k,k^{\prime}\in[K]$
$\widetilde{\operatorname{\operatorname{tr}}}_{k^{\prime},2}(\widetilde{P}_{k^{\prime}}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\tilde{\omega}_{k})\widetilde{P}_{k^{\prime}})=\delta_{k=k^{\prime}}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}.$
As a result,
$\widetilde{\mathcal{F}}(\mathcal{E}(x))=\sum_{k=1}^{K}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k}.$
It follows that
$\mathcal{D}(\widetilde{\mathcal{F}}(\mathcal{E}(x)))$ $=\sum_{k=1}^{K}U_{k}^{*}\widetilde{\operatorname{\operatorname{tr}}}_{k,2}(U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}\otimes\widetilde{\rho}_{k})U_{k}\otimes\rho_{k}$
$=\sum_{k=1}^{K}U_{k}^{*}U_{k}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})U_{k}^{*}U_{k}\otimes\rho_{k}$
$=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}=\mathcal{F}(x).$
We now show that the range and the shape vector of idempotent channels preserve the tensor product, which property will be used in the proofs of our Theorems 1 and 2.
Proposition 4.
For $\mathcal{F}$ an idempotent channel, the range of $\mathcal{F}^{*}$ and thus the shape of $\mathcal{F}$ satisfy for all $n\in\mathbb{N}^{*}$:
$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (13)
$\lambda(\mathcal{F}^{\otimes n})=\lambda(\mathcal{F})^{\otimes n}.$ (14)
The rest of this section is devoted to proving Propositions 1 and 4.
We begin by proving Proposition 1. We remark that the set of fixed points and the range of an idempotent function coincide.
Lemma 3.
$\operatorname{Rg}(f)\coloneqq\{f(x)~|~x\in\Omega\}=\{x\in\Omega~|~f(x)=x\}.$ (15)
Proof.
If $f(x)=x$, then $f(f(x))=f(x)=x$, so that $x\in\operatorname{Rg}(f)$. For the other inclusion, if $y=f(x)$, we also have $f(y)=f(f(x))=f(x)=y$, so $y$ is a fixed point. ∎
Proposition 1 basically follows from statements in [18] and we detail it here only for the convenience of the reader.
Proof of Proposition 1.
[…]
The structure of these algebras is well-understood (see for example [16, 5, 6]). In finite dimensions, any *-algebra can be decomposed as a direct sum of algebras of complex square matrices (which, in the theory of operator algebras, are called factors [16, 5, 6]). This allows us to write
$\operatorname{Rg}(\mathcal{F}^{*})\cong\bigoplus_{k=1}^{K}\mathcal{L}({\mathbb{C}}^{d_{k}}),$ (16)
for a set of integers $\{d_{1},\dots,d_{K}\}$, where $\cong$ denotes a *-algebra isomorphism. Furthermore, this decomposition is unique up to permutation of the terms in the right-hand side of eq. (16) and characterizes uniquely each *-algebra, in the sense that two *-algebras are isomorphic if and only if they have the same decomposition up to permutation of the terms in the direct sum.
We can go even further and give the general form of $\mathcal{F}^{*}$ using Proposition 1.5 in [18]. Indeed, we can decompose the Hilbert space $\mathcal{H}$ as a direct sum $\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, where we write $d_{k}=\dim(\mathcal{H}_{k,1})$ and $m_{k}=\dim(\mathcal{H}_{k,2})$, define a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ and a set of orthogonal projectors $\{P_{k}~|~\operatorname{Rg}(P_{k})=\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},k\in[K]\}$ such that $\mathcal{F}^{*}$ takes the form:
$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}}.$ (17)
The range $\operatorname{Rg}(\mathcal{F}^{*})$ will then be decomposed as
$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}.$ (18)
Indeed, let $x=\sum_{k=1}^{K}x_{k}\otimes\mathds{1}_{m_{k}}$ with $x_{k}\in\mathcal{L}(\mathcal{H}_{k,1})$ for all $k\in[K]$, then:
$\mathcal{F}^{*}(x)$ $=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(x_{k}\otimes\rho_{k})\otimes\mathds{1}_{m_{k}}=x,$
so that $x\in\operatorname{Rg}(\mathcal{F}^{*})$. This proves eq. (10). Note that eq. (18) implies eq. (16) as each term $\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}$ is isomorphic, as a *-algebra, to $\mathcal{L}(\mathcal{H}_{k,1})$.
Then, as $\mathcal{F}$ is the adjoint of $\mathcal{F}^{*}$, we can take the adjoint of eq. (17) and get:
$\forall x\in\mathcal{L}(\mathcal{H}),~\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}.$ (19)
Thus, whenever $\mathcal{F}$ is an idempotent channel, its range admits the decomposition
$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$ (20)
which proves eq. (11). ∎
We then proceed to the proof of Proposition 4, which will allow us to reduce the study of the problem of emulating $k$ copies of $\mathcal{F}$ using $n$ copies of $\mathcal{G}$ for general pairs of integers $(k,n)$ to the case where $k=n$. The fact that we can do such a reduction is probably the reason why we get a single-letter formula for the emulation capacity.
Proof of Proposition 4.
As, by definition, $\lambda(\mathcal{F})$ is obtained from the decomposition in factors of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$, it follows that eq. (14) is a consequence of eq. (13).
Let $n\in\mathbb{N}^{*}$, we take the $n$-th tensor power of eq. (17) to get that, for $x\in\mathcal{L}(\mathcal{H})$,
$\mathcal{F}^{*\otimes n}(x)=\sum_{\bar{k}\in[K]^{n}}\operatorname{\operatorname{tr}}_{\bar{k},2}(P_{\bar{k}}xP_{\bar{k}}(\mathds{1}_{d_{\bar{k}}}\otimes\rho_{\bar{k}}))\otimes\mathds{1}_{m_{\bar{k}}},$ (21)
with $\bar{k}=(k_{1},\dots,k_{n})$ ranging on all the $n$-uples in $K]^{n}$, $\operatorname{\operatorname{tr}}_{\bar{k},2}=\operatorname{\operatorname{tr}}_{k_{1},2}\otimes\dots\otimes\operatorname{\operatorname{tr}}_{k_{n},2}$ are the product partial traces, $P_{\bar{k}}=P_{k_{1}}\otimes\dots\otimes P_{k_{n}}$ are the product projections, $\rho_{\bar{k}}=\rho_{k_{1}}\otimes\dots\otimes\rho_{k_{n}}$ are product states, and similarly $\mathds{1}_{d_{\bar{k}}}=\mathds{1}_{d_{k_{1}}}\otimes\dots\otimes\mathds{1}_{d_{k_{n}}}$. Hence, the range of $\mathcal{F}^{*}$ can be decomposed as:
$\operatorname{Rg}(\mathcal{F}^{*\otimes n})=\bigoplus_{\bar{k}\in[K]^{n}}\mathcal{L}(\mathcal{H}_{\bar{k}})\otimes\mathds{1}_{m_{\bar{k}}}=\operatorname{Rg}(\mathcal{F}^{*})^{\otimes n},$ (22)
which completes the proof of Proposition 4. ∎
@@
Note that the shape of an idempotent channel $\mathcal{F}$ coincides with the shape of the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$. The notion of shape vector for *-algebras was introduced by Kuperberg in [11]. If $\mathcal{A}$ is a *-algebra, we can decompose
@@
for a set of integers $\{d_{k}~|~k\in[K]\}$ and we write $\lambda(\mathcal{A})$ the shape vector of $\mathcal{A}$, which coordinates are the integers $(d_{k})_{k\in[K]}$ sorted in non-increasing order. [^2] In fact, in all this article and in Kuperberg’s [11], the order with respect to which the coordinates of shape vectors are sorted is actually irrelevant.
4 Emulation of idempotent channel with zero error
@@
is achievable. In the second one, we prove that there is no rate strictly greater than eq. (23).
@@
Note that in the first case, i.e. $\lambda(\mathcal{F})$ has a coefficient greater or equal to two, the emulation capacity of $\mathcal{F}$ by the completely dephasing channel is zero, that is that it is impossible to emulate a channel preserving some quantum information with a classical one, without error.
@@
This example shows that neither the emulation capacity nor the strong converse rate given in Theorem 2 are reversible, since, in this example, the emulation capacity matches with the strong converse rate of Theorem 2, as $\mathcal{F}$ is taken to be an identity channel.
@@ 4.1 Achievability in Theorem 1
This theorem, despite not stated explicitly in [11], is proven in the proof of Theorem 1.1 of [11]. ∎
@@
Let $\mathcal{A}$, $\mathcal{B}$ be two finite-dimensional *-algebras and $\iota:\mathcal{A}\rightarrow\mathcal{B}$ be a subunital *-homomorphism. Let $\mathds{1}_{\mathcal{A}}\in\mathcal{A}$ and $\mathds{1}_{\mathcal{B}}\in\mathcal{B}$ be the identity operator of $\mathcal{A}$ and $\mathcal{B}$ respectively. Then, for all $x\in\mathcal{A}$,
$\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}})).$ (26)
Furthermore, $P_{\iota}\coloneqq\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector, and the map
$\widetilde{\iota}:x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ (27)
is unital completely positive. Finally, under the assumption that $\iota$ is injective, we write $\iota^{-1}:\iota(\mathcal{A})\subseteq\mathcal{B}\rightarrow\mathcal{A}$ its partial inverse, then
$\widetilde{\iota}^{-1}:x\in\mathcal{B}\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ (28)
is a unital completely positive map such that
$\forall x\in\mathcal{A},~\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)=x.$ (29)
Note that none of the two maps $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are *-homomorphisms unless $\iota$ is unital.
@@
We start by proving eq. (26). Let $x\in\mathcal{A}$, we can use the polarisation identity (eq. 1.45 in [18]) to write
$x=\frac{1}{4}\sum_{k=0}^{3}i^{k}(x+i^{k}\mathds{1}_{\mathcal{A}})^{*}(x+i^{k}\mathds{1}_{\mathcal{A}}),$
so that $x$ is decomposed as a linear combination of four positive operators in $\mathcal{A}$.
Thus, it suffices to show eq. (26) for positive semidefinite operators. For positive semidefinite $x\in\mathcal{A}$,
$0\leq x\leq\|x\|\mathds{1}_{\mathcal{A}}.$
Since $\iota$ is completely positive,
$0\leq\iota(x)\leq\|x\|\iota(\mathds{1}_{\mathcal{A}}).$
Thus, $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$. Now,
$\iota(\mathds{1}_{\mathcal{A}})^{2}=\iota(\mathds{1}_{\mathcal{A}}^{2})=\iota(\mathds{1}_{\mathcal{A}}),$
and, as $\iota$ is a *-homomorphism, it preserves the *-operation on $\mathcal{A}$, so that
$\iota(\mathds{1}_{\mathcal{A}})^{*}=\iota(\mathds{1}_{\mathcal{A}}).$

### v2

@@ 3 Decomposition properties of idempotent channels
The proof of eq. (19) and eq. (20) is elementary and these equations actually hold for every linear maps $\Phi:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{1}^{\prime})$ and $\Psi:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{H}_{2}^{\prime})$. Indeed, since $\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi)=\operatorname{\operatorname{Span}}\{y_{1}\otimes y_{2}~|~y_{1}\in\operatorname{Rg}(\Phi),y_{2}\in\operatorname{Rg}(\Psi)\}$, we argue on product elements $y_{1}\otimes y_{2}\in\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi)$ with $y_{1}\in\operatorname{Rg}(\Phi)$ and $y_{2}\in\operatorname{Rg}(\Psi)$. We have
$y_{1}\otimes y_{2}=\Phi(x_{1})\otimes\Psi(x_{2})=\Phi\otimes\Psi(x_{1}\otimes x_{2})\in\operatorname{Rg}(\Phi\otimes\Psi),$
where $x_{1}\in\mathcal{L}(\mathcal{H}_{1})$ and $x_{2}\in\mathcal{L}(\mathcal{H}_{2})$. Conversely, for $y\in\operatorname{Rg}(\Phi\otimes\Psi)$ we have $y=\Phi\otimes\Psi(x)$, with $x\in\mathcal{L}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})$. Thus, we can write $x$ as
$x=\sum_{j=1}^{J}x_{1}^{(j)}\otimes x_{2}^{(j)},$
with $x_{i}^{(j)}\in\mathcal{L}(\mathcal{H}_{i})$ for $(i,j)\in[2]\times[J]$. Therefore,
$y=\Phi\otimes\Psi(\sum_{j=1}^{J}x_{1}^{(j)}\otimes x_{2}^{(j)})=\sum_{j=1}^{J}\Phi(x_{1}^{(j)})\otimes\Psi(x_{2}^{(j)})\in\operatorname{Rg}(\Phi)\otimes\operatorname{Rg}(\Psi).$
Then, we prove the compatibility of taking the reduced version of idempotent channels with the tensor product. We write $\mathds{1}_{1}$ and $\mathds{1}_{2}$ respectively the identity operators in $\mathcal{L}(\mathcal{H}_{1})$ and $\mathcal{L}(\mathcal{H}_{2})$ and $\mathds{1}_{1,2}=\mathds{1}_{1}\otimes\mathds{1}_{2}$. We also write $\mathcal{H}_{i}^{(0)}=\operatorname{\operatorname{supp}}(\mathcal{F}_{i}(\mathds{1}_{i}))$ and $\mathcal{H}_{i}^{(1)}=\ker(\mathcal{F}_{i}(\mathds{1}_{i}))$ for $i\in[2]$ and the corresponding isometries $V_{i}:\mathcal{H}_{i}^{(0)}\rightarrow\mathcal{H}_{i}$ defining $\widehat{\mathcal{F}}_{i}$. Then, $\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})=\mathcal{F}_{1}(\mathds{1}_{1})\otimes\mathcal{F}_{2}(\mathds{1}_{2})$, hence $\operatorname{\operatorname{supp}}(\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2}))=\operatorname{\operatorname{supp}}(\mathcal{F}_{1}(\mathds{1}_{1}))\otimes\operatorname{\operatorname{supp}}(\mathcal{F}_{2}(\mathds{1}_{2}))$ and $\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$ maps $\mathcal{L}(\mathcal{H}_{1}^{(0)}\otimes\mathcal{H}_{2}^{(0)})$ into itself. If we write $e_{1,2}=\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})^{0}$ the orthogonal projector onto $\mathcal{H}_{1}^{(0)}\otimes\mathcal{H}_{2}^{(0)}$ the support of $\mathcal{F}_{1}\otimes\mathcal{F}_{2}(\mathds{1}_{1,2})$, we also have that $e_{1,2}=e_{1}\otimes e_{2}=(V_{1}\otimes V_{2})(V_{1}\otimes V_{2})^{*}$ with $e_{1}=V_{1}V_{1}^{*}=\mathcal{F}_{1}(\mathds{1}_{1})^{0}$ and $e_{2}=V_{2}V_{2}^{*}=\mathcal{F}_{2}(\mathds{1}_{2})^{0}$ the orthogonal projectors onto the respective supports of $\mathcal{F}_{i}(\mathds{1}_{i})$, $i\in[2]$. Then for $x_{1}\in\mathcal{L}(\mathcal{H}_{1})$ and $x_{2}\in\mathcal{L}(\mathcal{H}_{2})$, we have
$\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}(x_{1}\otimes x_{2})$ $=(V_{1}\otimes V_{2})^{*}\mathcal{F}_{1}\otimes\mathcal{F}_{2}((V_{1}\otimes V_{2})(x_{1}\otimes x_{2})(V_{1}\otimes V_{2})^{*})(V_{1}\otimes V_{2})$
$=V_{1}^{*}\mathcal{F}_{1}(V_{1}x_{1}V_{1}^{*})V_{1}\otimes V_{2}^{*}\mathcal{F}_{2}(V_{2}x_{2}V_{2}^{*})V_{2}$
$=\widehat{\mathcal{F}}_{1}(x_{1})\otimes\widehat{\mathcal{F}}_{2}(x_{2}).$
Eq. (21) then follows on the whole space $\mathcal{L}(\mathcal{H}_{1}\otimes\mathcal{H}_{2})=\mathcal{L}(\mathcal{H}_{1})\otimes\mathcal{L}(\mathcal{H}_{2})$ by linearity.
The last equality, eq. (22), is then seen by injecting eq. (8) into eq. (19) and by the definition of the shape vector of an idempotent channel. ∎
@@
Note that the shape of an idempotent channel $\mathcal{F}$ coincides with the shape of the *-algebra $\operatorname{Rg}(\widehat{\mathcal{F}}^{*})$, with $\widehat{\mathcal{F}}$ the reduced channel associated to $\mathcal{F}$. The notion of shape vector for *-algebras was introduced by Kuperberg in [12]. If $\mathcal{A}$ is a *-algebra, we can decompose
@@
for a set of integers $\{d_{k}~|~k\in[K]\}$ and we write $\lambda(\mathcal{A})$ the shape vector of $\mathcal{A}$, which coordinates are the integers $(d_{k})_{k\in[K]}$ sorted in non-increasing order. [^3] In fact, in all this article and in Kuperberg’s work [12], the order with respect to which the coordinates of shape vectors are sorted is actually irrelevant.
4 Emulation of idempotent channels with zero error
@@
is achievable. In the second one, we prove that any rate strictly greater than eq. (23) is not achievable.
@@
Note that in the first case, i.e. $\lambda(\mathcal{F})$ has a coefficient greater or equal to two, the emulation capacity of $\mathcal{F}$ by the completely dephasing channel is zero, that is it is impossible to emulate a channel preserving some quantum information with a classical one, without error.
@@
This example shows that neither the emulation capacity nor the strong converse rate given in Theorem 2 are reversible, since, as $\mathcal{F}$ is taken to be an identity channel in this example, both quantity match.
@@ 4.1 Achievability in Theorem 1
This theorem, despite not being stated explicitly in [12], is proven in the proof of Theorem 1.1 of [12]. ∎
@@
In the statement of the following lemma, we will use the fact that it is always possible, given $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ with the same identity operator as $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, to construct a unital completely positive map $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ such that for all $x\in\mathcal{A}$, $E_{\mathcal{A}}(x)=x$. This fact is easily seen as if $\mathcal{A}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ with the same identity operator, we can write the orthogonal direct sum decompositions
$\mathcal{H}_{\mathcal{A}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},\quad\mathcal{A}=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$
with $m_{k}=\dim(\mathcal{H}_{k,2})$ for $k\in[K]$. Then, we can take $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ defined by
$E_{\mathcal{A}}:x\rightarrow\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\frac{1}{m_{k}}\mathds{1}_{m_{k}},$ (26)
which is called the tracial conditional expectation on $\mathcal{A}$ (see e.g. Proposition 1.5 in [19]).
@@
Let $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, $\mathcal{B}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ be two finite-dimensional unital *-subalgebras of $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ respectively, where we assume that the identity operator in $\mathcal{A}$ coincides with the one in $\mathcal{L}(\mathcal{H}_{\mathcal{A}})$, denoted $\mathds{1}_{\mathcal{A}}$ and the one in $\mathcal{B}$ coincides with the one in $\mathcal{L}(\mathcal{H}_{\mathcal{B}})$, denoted $\mathds{1}_{\mathcal{B}}$. Let $\iota:\mathcal{A}\rightarrow\mathcal{B}$ be an injective subunital *-homomorphism.
Then, writing $\mathcal{H}_{\mathcal{B}}^{(0)}=\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))\subseteq\mathcal{H}_{\mathcal{B}}$, we have that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})\in\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ is the orthogonal projector on $\mathcal{H}_{\mathcal{B}}^{(0)}$. We denote $V:\mathcal{H}_{\mathcal{B}}^{(0)}\rightarrow\mathcal{H}_{\mathcal{B}}$ the isometry embedding $\mathcal{H}_{\mathcal{B}}^{(0)}$ into $\mathcal{H}_{\mathcal{B}}$ such that $VV^{*}=P_{\iota}$ and we define the reduced *-homomorphism $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ by
$\widehat{\iota}:x\mapsto V^{*}\iota(x)V.$ (27)
Then $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ is a unital injective *-homomorphism and $\widehat{\iota}(\mathcal{A})$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ where both $\widehat{\iota}(\mathcal{A})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ share the same identity operator. We will write $\widehat{\iota}^{-1}:\widehat{\iota}(\mathcal{A})\rightarrow\mathcal{A}$ the inverse of $\widehat{\iota}$, which is itself a unital injective *-homomorphism.
We write $E_{\mathcal{A}}:\mathcal{L}(\mathcal{H}_{\mathcal{A}})\rightarrow\mathcal{A}$ and $E_{\widehat{\iota}(\mathcal{A})}:\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})\rightarrow\widehat{\iota}(\mathcal{A})$ the tracial conditional expectations on $\mathcal{A}$ and $\widehat{\iota}(\mathcal{A})$ respectively. Then, the maps
[…]
$x$ $\mapsto\iota(E_{\mathcal{A}}(x))+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota}),$
and
$\widetilde{\iota}^{-1}:\mathcal{L}(\mathcal{H}_{\mathcal{B}})$ $\rightarrow\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{\mathcal{A}})$ (29)
$x$ $\mapsto\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(V^{*}xV))$
are unital completely positive and satisfy for all $x\in\mathcal{A}$,
$\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)=x.$ (30)
@@
First, we prove that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector on $\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$. As $\iota$ is a *-homomorphism, it preserves both the operation of taking the adjoint and the multiplication, so that $P_{\iota}^{*}=\iota(\mathds{1}_{\mathcal{A}})^{*}=\iota(\mathds{1}_{\mathcal{A}}^{*})=P_{\iota}$ and $P_{\iota}^{2}=\iota(\mathds{1}_{\mathcal{A}})^{2}=\iota(\mathds{1}_{\mathcal{A}}^{2})=P_{\iota}$.
We move on to prove that $\widehat{\iota}:\mathcal{A}\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$, with $\mathcal{H}_{\mathcal{B}}^{(0)}=\operatorname{\operatorname{supp}}(\iota(\mathds{1}_{\mathcal{A}}))$ is a *-homomorphism, which is also unital and injective. First $\widehat{\iota}$ is clearly linear. Let $x\in\mathcal{A}$,
$\widehat{\iota}(x)^{*}=(V^{*}\iota(x)V)^{*}=V\iota(x)^{*}V^{*}=V\iota(x^{*})V^{*}=\widehat{\iota}(x^{*}),$
and for $y\in\mathcal{A}$,
$\widehat{\iota}(xy)=V^{*}\iota(xy)V=V^{*}\iota(x)\iota(y)V\overset{(a)}{=}V^{*}VV^{*}\iota(x)VV^{*}\iota(y)V=\widehat{\iota}(x)\widehat{\iota}(y).$
The equality $(a)$ holds by applying Lemma 3 on the map $\Phi=\iota$ which is positive as it is a *-homomorphism [^4] Actually, *-homomorphisms are even completely positive [7].. To show the unitality, we have
$\widehat{\iota}(\mathds{1}_{\mathcal{A}})=V^{*}\iota(\mathds{1}_{\mathcal{A}})V=V^{*}P_{\iota}V=V^{*}VV^{*}V=\mathds{1}_{\mathcal{H}_{\mathcal{B}}^{(0)}},$
[…]
$0=\widehat{\iota}(x)=V^{*}\iota(x)V.$
Hence, by Lemma 3
$\iota(x)=VV^{*}\iota(x)VV^{*}=0,$
so that $x\in\ker(\iota)$, thus $x=0$ by injectivity of $\iota$.
Since $\widehat{\iota}$ is a unital *-homomorphism, $\widehat{\iota}(\mathcal{A})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$ and has the same identity operator as $\mathcal{L}(\mathcal{H}_{\mathcal{B}}^{(0)})$, so that the tracial conditional expectation $E_{\widehat{\iota}(\mathcal{A})}$ is well-defined by eq. (26). Furthermore, as $\widehat{\iota}$ is injective $\widehat{\iota}^{-1}$ is well-defined and is a unital *-homomorphism as well.
Now, we show that $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are unital completely positive maps. By unitality of $E_{\mathcal{A}}$, $E_{\widehat{\iota}(\mathcal{A})}$ and $\widehat{\iota}^{-1}$, it is clear that $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ are unital. Furthermore, $\widetilde{\iota}^{-1}$ is the composition of completely positive maps and is thus completely positive. Moreover, as $\iota$ was assumed to be subunital, $\iota(\mathds{1}_{\mathcal{A}})=P_{\iota}\leq\mathds{1}_{\mathcal{B}}$ so that $\widetilde{\iota}$ is completely positive.
We also have that $\widetilde{\iota}(\mathcal{L}(\mathcal{H}_{\mathcal{A}}))\subseteq\mathcal{B}$. Indeed, as $\mathcal{B}$ is unital of unit $\mathds{1}_{\mathcal{B}}$ and as the range of $\iota$ is in $\mathcal{B}$, $\mathds{1}_{\mathcal{B}}-P_{\iota}\in\mathcal{B}$ and $\iota(E_{\mathcal{A}}(x))\in\mathcal{B}$ for all $x\in\mathcal{L}(\mathcal{H}_{\mathcal{B}})$. Hence $\widetilde{\iota}(\mathcal{A})\subseteq\mathcal{B}$.